# FedOne: Query-Efficient Federated Learning for Black-box Discrete Prompt Learning

### v1

@@ A.4 Convergence Analysis of Federated Prompt Tuning
^$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})\stackrel{{\scriptstyle def}}{{=}}\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\hat{\nabla}_{\boldsymbol{\alpha}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})$ (92)
@@
Bound bias between local and average $\boldsymbol{\alpha}$. Let $\alpha_{i,j}\geq\nu>0$ for $i=1,...,n$ and $j=1,...,N$, $\eta$ is the learning rate, let $\mathbb{E}$ represents $\mathbb{E}_{\left\{\Phi^{k,r}_{t}\sim\text{GS}(\boldsymbol{\alpha}^{k}_{t})\right\}_{r=1}^{I}}$ and $\mathbb{E}_{\mathcal{B}_{t}^{k}}$, the bias between the local and the average $\boldsymbol{\alpha}$ can be bounded:
@@
$t_{c}\triangleq\lfloor\frac{t}{E}\rfloor E\$
$\boldsymbol{\alpha}_{i,\left(t_{c}\right)}=\frac{1}{K_{\ast}}\sum_{k\in U_{t_{c}}}\boldsymbol{\alpha}_{i,\left(t_{c}\right)}^{k}\$
@@
=(2)$\overset{(2)}{=}2\mathbb{E}\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\mathbb{E}\left[\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})\right]\right\|^{2}+2\left\|\mathbb{E}\left[\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})\right]\right\|^{2}$
@@
=$\overset{(3)}{=}2\mathbb{E}\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
=$\overset{(4)}{=}2\eta^{2}\sum_{\rho=t_{c}}^{t-1}\mathbb{E}\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
=$\overset{(5)}{=}2\eta^{2}\sum_{\rho=t_{c}}^{t-1}\mathbb{E}\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\eta^{2}(t-t_{c})\sum_{\rho=t_{c}}^{t-1}\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
where (1) use $10\leq\gamma\leq E-$ 1 and $1+...+E=\frac{E(E+1)}{2}$.
@@
Theorem 3.4. Suppose Assumption 3.1, 3.2 and 3.3 hold, for $t=0,1,...,T-1$, $B=\min\left\{B^{1},...,B^{K}\right\}$ where $B^{k}$ is the local mini-batch size. $\alpha_{i,j}\geq\nu>0$ for $i=1,...,n$ and $j=1,...,N$, $I$ is the sampling times for prompt. $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$ is the variance of the variance-reduced policy gradient, $\sigma_{\psi}^{2}$ is the variance of the stochastic gradient, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, and $\eta$ satisfies the following inequality:
@@
$\nabla_{\boldsymbol{\alpha}}\mathbb{E}_{\Phi_{t}\sim\text{GS}(\boldsymbol{\alpha}_{t})}\left[{\mathcal{L}}(\Phi_{t},\Psi)\right],\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\right\|^{2}$
@@
$\quad\ \mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\mathbb{E}_{U_{t}}\left[\left<\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})\right>\right]$
=$\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\mathbb{E}_{U_{t}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})\right]\right>$
=$\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{U_{t}}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\hat{\nabla}_{\boldsymbol{\alpha}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right>$
=$\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{U_{t}}\left[\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right>$
=$\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\frac{1}{K_{\ast}}\left[K_{\ast}\sum_{k=1}^{K}q^{\left[k\right]}\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right>$
=$\overset{(1)}{=}\frac{\eta}{2}\left[-\left\|\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k})\right\|^{2}-\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}+\left\|\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k})-\sum_{k=1}^{K}q^{\left[k\right]}\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right\|^{2}\right]$
@@
=$\overset{(1)}{=}\mathbb{E}_{U_{t}}\left\{\frac{L\eta^{2}}{2}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})-\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})\right]\right\|^{2}\right]\right\}$
@@
=$\overset{(2)}{=}\mathbb{E}_{U_{t}}\left\{\frac{L\eta^{2}}{2}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{\ast}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}\right]\right\}$
@@
=$\overset{1)}{=}c\left(c_{1}\sqrt{K_{\ast}}+c_{2}\frac{1}{\sqrt{K_{\ast}}}\right)^{2}$ (104)
@@ B.1 GLUE Dataset Metrics
Dataset $\mathbf{|L|}$ ||||Train|||| ||||Dev|||| ||||Test|||| Type Metrics Domain
@@ C.4 Test Metrics for the RoBERTa-Large Experiment for FedAvg
Fed-PromptTuning ${41.8}_{0.0}$ ${66.3}_{0.1}$ ${78.3}_{0.2}$ ${80.7}_{0.2}$ ${0.6}_{0.0}$ ${51.1}_{0.0}$ ${51.4}_{0.3}$ 52.8952.89{52.89}52.89
Fed-P-Tuning v2 ${42.4}_{0.1}$ ${66.6}_{0.1}$ ${82.4}_{0.5}$ ${80.6}_{0.4}$ ${0.9}_{0.1}$ ${52.9}_{0.1}$ ${53.7}_{0.4}$ 54.2154.21{54.21}54.21
Fed-BBT ${41.4}_{1.2}$ ${66.7}_{0.2}$ ${77.6}_{0.4}$ ${80.1}_{1.2}$ ${2.7}_{0.6}$ ${51.9}_{0.6}$ ${52.3}_{0.7}$ 53.2453.24{53.24}53.24
Fed-BDPL ${40.8}_{1.0}$ ${67.0}_{0.2}$ ${82.9}_{2.8}$ ${80.5}_{1.6}$ ${5.9}_{0.2}$ ${52.0}_{0.6}$ ${53.8}_{0.5}$ 54.7054.70{54.70}54.70
Fed-GS-BDPL ${41.2}_{0.4}$ ${67.1}_{0.1}$ ${82.8}_{1.4}$ ${80.5}_{1.0}$ ${5.8}_{0.5}$ ${52.3}_{0.2}$ ${54.5}_{0.4}$ 54.8954.89{54.89}54.89

### v2

@@ A.4 Convergence Analysis of Federated Prompt Tuning
$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})\stackrel{{\scriptstyle def}}{{=}}\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\hat{\nabla}_{\boldsymbol{\alpha}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})$ (92)
@@
Bound bias between local and average $\boldsymbol{\alpha}$. Let $\alpha_{i,j}\geq\nu>0$ for i=1,,ni=1,...,n and j=1,,Nj=1,...,N, $\eta$ is the learning rate, let $\mathbb{E}$ represents $\mathbb{E}_{\left\{\Phi^{k,r}_{t}\sim\text{GS}(\boldsymbol{\alpha}^{k}_{t})\right\}_{r=1}^{I}}$ and $\mathbb{E}_{\mathcal{B}_{t}^{k}}$, the bias between the local and the average $\boldsymbol{\alpha}$ can be bounded:
@@
$t_{c}\triangleq\lfloor\frac{t}{E}\rfloor E\\$
$\boldsymbol{\alpha}_{i,\left(t_{c}\right)}=\frac{1}{K_{\ast}}\sum_{k\in U_{t_{c}}}\boldsymbol{\alpha}_{i,\left(t_{c}\right)}^{k}\\$
@@
$\overset{(2)}{=}2\mathbb{E}\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\mathbb{E}\left[\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})\right]\right\|^{2}+2\left\|\mathbb{E}\left[\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})\right]\right\|^{2}$
@@
$\overset{(3)}{=}2\mathbb{E}\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
$\overset{(4)}{=}2\eta^{2}\sum_{\rho=t_{c}}^{t-1}\mathbb{E}\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\left\|\sum_{\rho=t_{c}}^{t-1}\eta\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
$\overset{(5)}{=}2\eta^{2}\sum_{\rho=t_{c}}^{t-1}\mathbb{E}\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{\rho},\mathcal{B}^{k}_{\rho})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}+2\eta^{2}(t-t_{c})\sum_{\rho=t_{c}}^{t-1}\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{\rho},\Psi^{k})\right\|^{2}$
@@
where (1) use $0\leq\gamma\leq E-1$ and $1+...+E=\frac{E(E+1)}{2}$.
@@
Theorem 3.4. Suppose Assumption 3.1, 3.2 and 3.3 hold, for t=0,1,,T1t=0,1,...,T-1, $B=\min\left\{B^{1},...,B^{K}\right\}$ where $B^{k}$ is the local mini-batch size. $\alpha_{i,j}\geq\nu>0$ for i=1,,ni=1,...,n and j=1,,Nj=1,...,N, II is the sampling times for prompt. $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$ is the variance of the variance-reduced policy gradient, $\sigma_{\psi}^{2}$ is the variance of the stochastic gradient, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, and $\eta$ satisfies the following inequality:
@@
$\leq\left<\nabla_{\boldsymbol{\alpha}}\mathbb{E}_{\Phi_{t}\sim\text{GS}(\boldsymbol{\alpha}_{t})}\left[{\mathcal{L}}(\Phi_{t},\Psi)\right],\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\right\|^{2}$
@@
[…]
$=\left<\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\mathbb{E}_{U_{t}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})\right]\right$>
$=\left<\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{U_{t}}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\hat{\nabla}_{\boldsymbol{\alpha}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right$>
$=\left<\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{U_{t}}\left[\frac{1}{K_{\ast}}\sum_{k\in U_{t}}\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right$>
$=\left<\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k}),-\eta\cdot\frac{1}{K_{\ast}}\left[K_{\ast}\sum_{k=1}^{K}q^{\left[k\right]}\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right]\right$>
$\overset{(1)}{=}\frac{\eta}{2}\left[-\left\|\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k})\right\|^{2}-\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}+\left\|\nabla_{\boldsymbol{\alpha}_{i}}F(\boldsymbol{\alpha}_{t},\Psi^{k})-\sum_{k=1}^{K}q^{\left[k\right]}\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})\right\|^{2}\right]$
@@
$\overset{(1)}{=}\mathbb{E}_{U_{t}}\left\{\frac{L\eta^{2}}{2}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})-\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})\right]\right\|^{2}\right]\right\}$
@@
$\overset{(2)}{=}\mathbb{E}_{U_{t}}\left\{\frac{L\eta^{2}}{2}\mathbb{E}_{\left\{\Phi^{k,r}\right\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}^{k}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\boldsymbol{\alpha}_{t}^{k},\mathcal{B}_{t}^{k})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{\ast}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}\right]\right\}$
@@
$\overset{1)}{=}c\left(c_{1}\sqrt{K_{\ast}}+c_{2}\frac{1}{\sqrt{K_{\ast}}}\right)^{2}$ (104)
@@ B.1 GLUE Dataset Metrics
Dataset $\mathbf{|L|}$ ||Train|| ||Dev|| ||Test|| Type Metrics Domain
@@ C.4 Test Metrics for the RoBERTa-Large Experiment for FedAvg
Fed-PromptTuning ${41.8}_{0.0}$ ${66.3}_{0.1}$ ${78.3}_{0.2}$ ${80.7}_{0.2}$ ${0.6}_{0.0}$ ${51.1}_{0.0}$ ${51.4}_{0.3}$ 52.89{52.89}
Fed-P-Tuning v2 ${42.4}_{0.1}$ ${66.6}_{0.1}$ ${82.4}_{0.5}$ ${80.6}_{0.4}$ ${0.9}_{0.1}$ ${52.9}_{0.1}$ ${53.7}_{0.4}$ 54.21{54.21}
Fed-BBT ${41.4}_{1.2}$ ${66.7}_{0.2}$ ${77.6}_{0.4}$ ${80.1}_{1.2}$ ${2.7}_{0.6}$ ${51.9}_{0.6}$ ${52.3}_{0.7}$ 53.24{53.24}
Fed-BDPL ${40.8}_{1.0}$ ${67.0}_{0.2}$ ${82.9}_{2.8}$ ${80.5}_{1.6}$ ${5.9}_{0.2}$ ${52.0}_{0.6}$ ${53.8}_{0.5}$ 54.70{54.70}
Fed-GS-BDPL ${41.2}_{0.4}$ ${67.1}_{0.1}$ ${82.8}_{1.4}$ ${80.5}_{1.0}$ ${5.8}_{0.5}$ ${52.3}_{0.2}$ ${54.5}_{0.4}$ 54.89{54.89}